Provide ALL possible paths for the completion of the PROJECT 1. Calculate total time required to complete each path. Determin Question on Project 2, Steps. Help. Thanks
Answer #1

Project 2 * Network diagram. * Forward pass (Earliest times). set 1, -0 0₂= D12=1 Az = 02 + D23 = 1+2=3 Duran Dz+D34, O₂ + D2Date As = min 2 46-D56 A-D57? ord = minf g - 2 9 -37 svody no As = 6 AL= A5-Dus = 6-3=3 A 2 to Az A4-D34=3 0 =3 A = min 24₂-D